Consider three random events, A, B and C. Suppose that P(A) = 0.5, P(A∩C) = 0.2, P(C) = 0.4, P(B) = 0.4, P(A∩B∩C) = 0.1, P(B∩C) = 0.18, and P(A∩B) = 0.21. Calculate the following probabilities:

c. P((B∩C)c ∪(A∩B)c)
Answer #1

solution: P(A)=0.5,P(B)=0.4,P(C)=0.4

P(A C) = 0.2,

P(B\cap C)=0.18

P(A\cap B)=0.21

P(A\cap B\cap C)=0.1

we have to calculate P((B\cap C)^{c}\cup (A\cap B)^{c})

P(B\cap C)^{c}=1-P(B\cap C)=1-0.18=0.82

P(A\cap B)^{c}=1-P(A\cap B)=1-0.21=0.79

P((B\cap C)^{c}\cap (A\cap B)^{c})=1-((A\cap B)+(B\cap C)-(A\cap B\cap C))=1-(0.21+0.18-0.1)=1-0.29=0.71

P((B\cap C)^{c}\cup (A\cap B)^{c})=P(B\cap C)^{c}+P(A\cap C)^{c}-P((B\cap C)^{c}\cap (A\cap B)^{c})

P((B\cap C)^{c}\cup (A\cap B)^{c})=0.82+0.79-0.71=0.9